OpenStudy (anonymous):

What is the relative maximum and minimum of the function? f(x) = 2x2 + 28x - 8

OpenStudy (anonymous):

x=-7

OpenStudy (anonymous):

A. Minimum Value: -106 Range y > -106 B. Minimum Value: -7 Range y > -7 C. Minimum Value: 7 Range y > 7 D. Minimum Value: -106 Range y > -7

OpenStudy (anonymous):

is it B

OpenStudy (ivancsc1996):

Well, since this is a parabola going upwards you just need to equal the dervative to zero and solve for x.\[\frac{ df }{ dx }=4x+28=0 \rightarrow x=\frac{ -28 }{ 4 }=-7\] Then solve for f.\[f=114\]Thats the minimum. The maximum is infinity because as x goes to either side of minus seven, f gets bigger and bigger.

OpenStudy (homeworkhelper099):

That answer is wrong